OpenStudy (anonymous):

how would i simplify this complex fraction? 1/x+1 divided by -1/x+1

OpenStudy (thesecret20111):

Are you asking \[\frac{ 1 }{ x }+\frac{ \frac{ -1 }{ 1 } }{ x }+1\] ?

OpenStudy (thesecret20111):

If so the\[\frac{ \frac{ -1 }{ 1 } }{ x }\] part equals -1/x

OpenStudy (thesecret20111):

Therefore you would have 1/x + -1/x + 1 1/x + -1/x =0

OpenStudy (thesecret20111):

Therefore you would have 0+1 which equals 1 :)